A rectangle has an area of 96 square feet. If the width of the rectangle is six times the length, what is the width in feet?

Answer by RAY100(1637)

Area = Length * Width
.
A=96 = L*W,,,,,but W=6*L,,,,given
,
96 = L*6L = 6L^2
.
16 = L^2
.
L= +/- 4,,,,,,use +4
.
W=6*4 = 24
.
.
check,,,,,,4+24 = 96,,,,,ok
.